# Syntomic cohomology of truncated Brown–Peterson spectra

### v1

@@ 3. Hodge Tate Cohomology
On bigraded homotopy groups, the commutative diagram
@@
First, note that there is a canonical map
@@
This produces a element $\mu^{p^{n+1}}$ in the bigraded commutative commutative ring
@@
as a map of $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n}}]$-algebras and it makes sense to ask whether the map is given by inverting the class $\mu^{p^{n}}$. As in [HW22, Recollection 4.3.2], we can determine that the associated graded $\mathrm{gr}_{\textup{Ad}}^{*}\operatorname{BP}\langle n\rangle$ of the filtration
@@
and in fact the same considerations as in the proof of [HW22, Theorem 4.3.1] allow us to determine that the map is given by inverting $\mu$ on the $\operatorname{\mathbb{E}}_{2}$-pages. We have completely computed the top spectral sequence and since the bottom spectral sequence is given by inverting $\mu$ on $\operatorname{\mathbb{E}}_{2}$-pages, we determine that at each page the spectral sequence is given by inverting some power of $\mu$, and hence remains a monomomorphism entirely determining also the spectral sequence on the bottom. This proves the claim. ∎
Theorem 3.2.
@@
and converging to $\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$ collapses after the $d_{1}$ differential. All permanent cycles are on the 0-line.
@@
Since $v_{n+1}$ is detected by $t\mu^{p^{n+1}}$, we determine that $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$. Using the action of $\lambda_{1},\cdots,\lambda_{n+1}$ together with the monomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\cdots,\lambda_{n+1})$, we determine a Leibniz rule $d_{1}(x\varepsilon_{n+1})=xt\mu^{p^{n+1}}$ for all
@@
Since $\mu$ is a unit in a ring that acts on this spectral sequence, mainly the corresponding spectral sequence for $\tau_{\geq 0}L_{n}^{f}\operatorname{MU}$, the spectral sequence collapses after running the first differential and all permanent cycles are concentrated on the 0-line. Moreover, we can conclude that
@@ 4. Prismatic cohomology
of $\operatorname{\mathbb{E}}_{0}$-$\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP})$-algebras induced by the canonical $\operatorname{\mathbb{E}}_{1}$-$\operatorname{BP}$-algebra reduction map
@@
for $\ell\in\mathbb{Z}$, $0<j<p$ and $S,T\subset\{1,\cdots,n+1\}$ with $m\not\in S$. The spectral sequence then collapses at the $p^{n+1}+1$-page without room for further differentials. Consequently,
@@
The differential $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$ follows from the definition of $\varepsilon_{n+1}$ and Theorem 2.15. The Leibniz rule
@@
for all $x\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})[t^{\pm 1}]$ follows from [AHW24b, Corollary 3.2.6], which implies that the classes $\lambda_{1},\cdots,\lambda_{n+1}$ are permanent cycles that act on this spectral sequence, together with the monomomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\cdots,\lambda_{n+1})$.
@@
for all $S\subset\{1,\cdots,n+1\}$ with $m\not\in S$ for each $m\geq 1$ follow from the action of the spectral sequence for $\operatorname{BP}$ and the differentials in that spectral sequence, computed in [AHW24b, Proposition 3.2.9]. Finally, we simply check bidegrees and observe that the spectral sequence must collapse at the $\mathrm{E}_{p^{n+1}+1}$-page. ∎
@@ 5. Syntomic cohomology
We note that these bigraded groups are equipped with the Nygaard filtration coming the $t$-Bockstein spectral sequence and therefore there is an exact sequence
@@
$A_{10}$ $:=\mathrm{Nyg}_{\geq 1}\cap T\,$,
$A_{00}$ $:=\mathrm{Nyg}_{\geq 1}/A_{10}\,$,
$A_{11}$ $:=T/A_{10}\text{ and }$
$A_{10}$ $=F/A_{00}\,.$
Property 5.2.
@@
$A_{10}:=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{t^{dp^{j-1}}\lambda_{j}:0<d<p\}\,.$
@@
The following corollary is immediate.
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{p^{n+1}}]\oplus$
$\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{d,j}:0<d<p\}$
as a $\mathbb{F}_{p}\langle\lambda_{1},\cdots\lambda_{n+1}\rangle$-module where $\Xi_{d,j}$ is detected by $t^{dp^{j-1}}\lambda_{j}$, $y^{p^{n+1}}$ is detected by $t^{-p^{n+1}}$ and we abuse notation and write $\lambda_{k}$ for the class detected by $\lambda_{k}$. Moreover, the canonical map
@@
sends $\lambda_{k}$ to $\lambda_{k}$, $y^{p^{n+1}}$ to $y^{p^{n+1}}$, and sends $\lambda_{s}^{\epsilon}\cdot\Xi_{d,j}$ to zero for each $\epsilon\in\{0,1\}$, $s\in\{1,\cdots,n\}-\{j\}$, and $j\in\{1,\cdots,n+1\}$.
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,\cdots,v_{n+1})[t]\implies\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\otimes\mathbb{F}[\mu^{\pm p^{n}},t]\langle\varepsilon_{n+1}\rangle$
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{\pm p^{n+1}}]$
where $y^{p^{n+1}}$ is detected by $\mu^{p^{n+1}}$.
Letting
$A$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
$B$ $=\bigoplus_{j=1}^{n+1}\bigotimes_{s\in\{1,\cdots,n+1\}-\{j\}}\mathbb{F}_{p}\langle\lambda_{s}\rangle\{\Xi_{j,d}:0<d<p\}\,,\text{ and }$
$C$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{p^{n+1}}]\{y^{p^{n+1}}\}$
@@
is given by
$\varphi|_{A}=\textup{inc}_{A}\circ\mathrm{id}_{A}$
and $\varphi|_{B\oplus C}=0$ where $\textup{inc}_{A}:A\to A\oplus B\oplus C$ is the canonical inclusion.
@@
The mod $(p,\dots,v_{n+1})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
(5.6) $\mathbb{F}_{p}\langle\delta,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
@@
$A$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
$B$ $=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}\,$,
$C$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{-p^{n+1}}]\{y^{-p^{n+1}}\}$
then we determined that $\mathrm{can}|_{A}=\textup{incl}_{A\oplus C}\circ\mathrm{id}_{A\oplus C}$ and $\mathrm{can}(B)=0$ in Corollary 5.3, where $\textup{incl}_{A\oplus C}$ denotes the canonical inclusion $A\oplus C\to A\oplus B\oplus C$. We determined that $\varphi|_{A}=\textup{inc}_{A}\circ\mathrm{id}_{A}$ and $\varphi|_{B\oplus C}=0$ in Proposition 5.4. We therefore determine that $\ker(\mathrm{can}-\varphi)=A\oplus B$ and $\operatorname{coker}(\mathrm{can}-\varphi)=A$. We write $\partial$ to indicate that these classes come from the boundary map in the long exact sequence mod $(p,v_{1},\cdots,v_{n+1})$ induced by the fiber sequence
@@
The mod $(p,\dots,v_{n})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
(5.8) $\mathbb{F}_{p}[v_{n+1}]\langle\delta,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}[v_{n+1}]\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}.$
@@
for all primes $p>2$. This handles all the $\mathbb{F}_{p}[v_{n+1}]\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$-module generators and therefore proves the claim and therefore the $v_{n+1}$, $\partial$, $\lambda_{1}$, $\cdots$ $\lambda_{n+1}$-linearity proves the claim. ∎
Remark 5.9.
The dimension of the syntomic cohomology of $\operatorname{BP}\langle n\rangle$ mod $(p,v_{1},\cdots,v_{n+1})$ was predicted by Ausoni–Rognes [AR02, p. 7], as stated precisely in [ROG01, §2]. They also claim that when $\operatorname{BP}\langle n\rangle$ is an $\operatorname{\mathbb{E}}_{\infty}$-ring and a Smith–Toda complex $\operatorname{\mathbb{S}}/(p,v_{1},\cdots,v_{n})$ exists, then the $v_{n+1}$-Bockstein spectral sequence and the motivic spectral sequence collapse. It follows from [AAC+25], that at least when $p\geq 7$ and $n=2$ it suffices to know that $\operatorname{BP}\langle n\rangle$ is an $\operatorname{\mathbb{E}}_{3}$-ring, which is known by [HW22].

### v2

@@ 3. Hodge Tate Cohomology
The commutative diagram
@@
First, note that by construction there is a map
@@
This produces an element $\mu^{p^{n+1}}$ in the bigraded commutative commutative ring
@@
as a map of $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n}}]$-modules and it makes sense to ask whether the map is given by inverting the class $\mu^{p^{n}}$. As in [HW22, Recollection 4.3.2], we can determine that the associated graded $\mathrm{gr}_{\textup{Ad}}^{*}\operatorname{BP}\langle n\rangle$ of the filtration
@@
and in fact the same considerations as in the proof of [HW22, Theorem 4.3.1] allow us to determine that the map is given by inverting $\mu$ on the $\mathrm{E}_{2}$-pages. We have completely computed the top spectral sequence and since the bottom spectral sequence is given by inverting $\mu$ on $\mathrm{E}_{2}$-pages, we determine that at each page the spectral sequence is given by inverting some power of $\mu$, and hence remains a monomomorphism entirely determining also the spectral sequence on the bottom. This proves the claim. ∎
Proposition 3.2.
@@
and converging to $\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$ (up to isomorphism) collapses after the $d_{1}$ differential. All permanent cycles are on the 0-line.
@@
Since $v_{n+1}$ is detected by $t\mu^{p^{n+1}}$, we determine that $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$ by Definition 2.16 ($\varepsilon_{n+1}$ corresponds to a null homotopy of $v_{n+1}$). Using the action of $\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}$ together with the monomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1})$, we determine a Leibniz rule $d_{1}(x\varepsilon_{n+1})=xt\mu^{p^{n+1}}$ for all
@@
Since $\mu$ is a unit in an algebra spectral sequence that acts on this spectral sequence, namely the corresponding spectral sequence for $\tau_{\geq 0}L_{n}^{f}\operatorname{MU}$ (cf. proof of Theorem 3.1), the spectral sequence collapses after running the first differential and all permanent cycles are concentrated on the 0-line. Moreover, we can conclude that
@@ 4. Prismatic cohomology
of $\operatorname{\mathbb{E}}_{0}$ $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP})$-algebras induced by the canonical $\operatorname{\mathbb{E}}_{1}$ $\operatorname{BP}$-algebra reduction map
@@
for $0<j<p$ and $S\subset\{1,\cdots,n+1\}-\{m\}$. The spectral sequence then collapses at the $p^{n+1}+1$-page without room for further differentials. Consequently,
@@
The differential $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$ follows from Definition 2.16 ($\varepsilon_{n+1}$ corresponds to a null homotopy of $v_{n+1}$) and Theorem 2.17. The Leibniz rule
@@
for all $x\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})[t^{\pm 1}]$ follows from [AHW24b, Corollary 3.2.6], which implies that the classes $\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}$ are permanent cycles that act on this spectral sequence, together with the monomomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1})$.
@@
for all $S\subset\{1,\cdots,n+1\}$ with $m\not\in S$ for each $m\geq 1$ follow from the action of the spectral sequence for $\operatorname{BP}$ and the differentials in that spectral sequence, computed in [AHW24b, Proposition 3.2.9]. By [AHW24b, Proposition 3.2.9], we also know that that $t^{p^{n+1}}$ survives to the $\mathrm{E}_{p^{n+2}}$-page and it acts on the spectral sequence up until this page, so $\mathrm{E}_{p^{n+1}+1}=\mathrm{E}_{p^{n+2}}$. Finally, we simply check bidegrees and observe that the spectral sequence must collapse at the $\mathrm{E}_{p^{n+2}}$-page. ∎
Remark 4.1.
Note that, as a consequence of Proposition 4.2, we have an isomorphism
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\cong\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,v_{1},\cdots,v_{n})\,.$
This could alternatively be proven directly along the lines of [HRW22, § 6.4] and we believe that such a result holds quite generally, however one still needs to understand the Nygaard filtration on prismatic cohomology to compute syntomic cohomology and this is also achieved in Proposition 4.2
@@ 5. Syntomic cohomology
We note that these bigraded groups are equipped with the Nygaard filtration coming from the $t$-Bockstein spectral sequence and therefore there is an exact sequence
@@
$A_{11}$ $:=\mathrm{Nyg}_{\geq 1}\cap T\,$,
$A_{10}$ $:=\mathrm{Nyg}_{\geq 1}/A_{11}\,$,
$A_{01}$ $:=T/A_{11}\text{ and }$
$A_{00}$ $=F/A_{10}\,.$
Proposition 5.2.
@@
$A_{11}=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{t^{dp^{j-1}}\lambda_{j}:0<d<p\}\,.$
@@
The following corollary is immediate from Proposition 4.2 and Proposition 5.2.
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})=A_{00}\oplus A_{10}\oplus A_{01}\oplus A_{11}$
as $\mathbb{F}_{p}\langle\lambda_{1},\lambda_{2},\cdots\lambda_{n+1}\rangle$-modules where
$A_{00}$ $=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
$A_{11}$ $=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{\Xi_{d,j}:0<d<p\}\,$,
$A_{01}$ $=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}]\{\mu^{p^{n+1}}\}\text{ and }$
$A_{10}$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[t^{p^{n+1}}]\{t^{p^{n+1}}\}\,.$
Here $\Xi_{d,j}$ is detected by $t^{dp^{j-1}}\lambda_{j}$, we abuse notation and write $t^{p^{n+1}}$ for the class detected by $t^{p^{n+1}}$, we write $\lambda_{k}$ for the class detected by $\lambda_{k}$ and we write $\mu^{kp^{n+1}}$ for the class detected by $\mu^{kp^{n+1}}$. Moreover, the canonical map
@@
sends $\lambda_{k}$ to $\lambda_{k}$, $t^{p^{n+1}}$ to $t^{p^{n+1}}$, and sends $\lambda_{S}\cdot\Xi_{d,j}$ and $\lambda_{T}\mu^{kp^{n+1}}$ to zero for each $k\geq 1$, $S\subset\{1,\cdots,n+1\}-\{j\}$, $j\in\{1,\cdots,n+1\}$, $0<d<p$ and $T\subset\{1,\cdots,n+1\}$. In other words, $\mathrm{can}|_{A_{01}\oplus A_{11}}=0$ and $\mathrm{can}|_{A_{00}\oplus A_{10}}=\mathrm{inc}_{A_{00}\oplus A_{10}}$ where $\textup{inc}_{A_{00}\oplus A_{10}}$ is the canonical inclusion.
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,\cdots,v_{n+1})[t]\implies\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}(\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}})^{hS^{1}}/(p,v_{1},\cdots,v_{n+1})$
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{\pm p^{n}},t]\langle\varepsilon_{n+1}\rangle$
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}(\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}})^{hS^{1}}/(p,v_{1},\cdots,v_{n+1})=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{\pm p^{n+1}}]$
where $y^{\pm kp^{n+1}}$ is detected by $\mu^{\pm kp^{n+1}}$.
@@
is given by $\varphi|_{A_{00}}=\textup{inc}_{A_{00}}$, $\varphi|_{A_{01}}$ is an isomorphism and $\varphi|_{A_{10}\oplus A_{11}}=0$, where $\textup{inc}_{A_{00}}$ is the canonical inclusion.
Proof.
This follows from Proposition 3.2 and Corollary 5.3 together with [AHW24b, Corollary 3.2.6], which implies that the classes in $\mathbb{F}_{p}\langle\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}\rangle$ are equalized by $\mathrm{can}$ and $\varphi$. ∎
@@
Let $n\geq-1$. The mod $(p,v_{1},\dots,v_{n+1})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
(5.6) $\mathbb{F}_{p}\langle\partial,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
@@
$A_{00}$ $=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
$A_{10}$ $=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{t^{dp^{j-1}}\lambda_{j}:0<d<p\}\,$,
$A_{01}$ $=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}]\{\mu^{p^{n+1}}\}\text{ and }$
$A_{11}$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[t^{p^{n+1}}]\{t^{p^{n+1}}\}\,.$
We determined that $\mathrm{can}|_{A_{01}\oplus A_{11}}=0$ and $\mathrm{can}|_{A_{00}\oplus A_{10}}=\mathrm{inc}_{A_{00}\oplus A_{10}}$ in Corollary 5.3. We determined that $\varphi|_{A_{00}}=\textup{inc}_{A_{00}}$, $\varphi|_{A_{01}}$ is an isomorphism and $\varphi|_{A_{10}\oplus A_{11}}=0$ in Proposition 5.4. We therefore conclude that $\ker(\mathrm{can}-\varphi)=A_{00}\oplus A_{11}$ and $\operatorname{coker}(\mathrm{can}-\varphi)=A_{00}$. We write $\partial$ to indicate that classes that are multiples $\partial$ of come from the boundary map in the long exact sequence mod $(p,v_{1},\cdots,v_{n+1})$ induced by the fiber sequence
@@
Let $n\geq-1$. The mod $(p,\dots,v_{n})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
(5.8) $\mathbb{F}_{p}[v_{n+1}]\langle\partial,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}[v_{n+1}]\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}.$
@@
for all primes $p>2$. This handles all the $\mathbb{F}_{p}[v_{n+1}]\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$-module generators and therefore the $v_{n+1}$, $\partial$, $\lambda_{1}$, $\cdots$ $\lambda_{n+1}$-linearity proves the claim. ∎